Hi there,
I am having trouble with finding an equation for two sets of coordinates.
for example, the question is
find the equation for the coordinates (0,-3) (2,3)
I know y=mx+c and I know how to find m and c but I don't understand some things.
would they have the same equation or not? because for x1 y1 the equation would be
y=3x-3 right? But I don't think it is right for the whole of the coordinates because shouldn't there only be one equation for the co-ordinate (0,-3) (2,3)?
Help or if you cant understand me point me in the direction of a page to help :)
Thanks
Sariah
p.s this site is amazing :) helped out a lot.

Answer by lwsshak3(11628)

find the equation for the coordinates (0,-3) (2,3)
I think you are referring to the equation of a line, y=mx+b, m=slope, b=y-intercept
slope can be found when you have 2 points on the line which you have.
slope=∆y/∆x=(y1-y2)/(x1-x2)=(3-(-3))/(2-0)=6/2=3
equation: y=3x+b
solve for b, using any one of the given points(2,3)
3=3*2+b
b=-3
equation of the line given coordinates are on: y=3x-3
You got the correct answer and hopefully you now have a better understanding of the problem.